OpenStudy (anonymous):

An object has a mass of 12kg. On Planet A, the object weighs 117.6 N. Is the force of gravity on Planet A equal to, less than or greater than the force of gravity on Earth?

OpenStudy (anonymous):

Gravity on Earth:g=9.8m/s^2 Using W=m*g W(on earth)=12*9.8=117.6 N So the force of gravity is the same.

OpenStudy (anonymous):

Makes sense?

OpenStudy (anonymous):

Yes, starting to. Thanks.

OpenStudy (anonymous):

You are welcome! (Thanks for the "best response") ;)